OpenStudy (anonymous):

i need help with the square root property x^2=12

OpenStudy (raden):

so, x = sqrt(12), right ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

oppss, i forgot give sign +- before squareroot it should be : 

OpenStudy (raden):

then simplify : sqrt(12) = sqrt(4*3) = sqrt(4) * sqrt(3) = 2 sqrt(3)